Going into the final exam, which counts as two grades, you have the following test scores of 84, 78, 64, and 88. What score do you need on the final exam in order to have an average of 80?
Answer by nerdybill(7384)

Going into the final exam, which counts as two grades, you have the following test scores of 84, 78, 64, and 88. What score do you need on the final exam in order to have an average of 80?
Let x = final exam score
the
(x + x + 84 + 78 + 64 + 88)/6 = 80
(2x + 84 + 78 + 64 + 88)/6 = 80
(2x + 314)/6 = 80
2x + 314 = 480
2x = 166

x = 83